# Do you monitor CI Practices? I don’t know. You tell me: A case study

### v1

@@ 3.1. Survey
To gain an understanding of the project’s CI maturity before implementing the case study, we requested each developer to answer a preliminary survey. This initial survey aided in comprehending the CI state within each project prior to the case study, assessing satisfaction with CI adoption, and identifying any metrics used to measure CI. This preliminary insight facilitated a more comprehensive evaluation of the case study’s impact on the evolution of CI in each analyzed project. Furthermore, the survey responses helped us answer RQ1.
@@
The survey consisted of 15 questions, 9 of which were open-ended. To give respondents as much flexibility as possible, 9 questions were mandatory (6 of which were close-ended while 3 was open-ended), and 6 questions were optional (all open-ended). The survey was divided into four sections: 1) Introduction, 2) Personal information, 3) Team Organization And Collaboration, and 4) Diagnosis of the project’s CI maturity. A full copy of the survey is available at https://zenodo.org/records/11106110.
@@
In the “Introduction” section of the survey, we provide context for the study, introduce Continuous Integration concept, and outline the survey’s objectives. We also specify the estimated time required for completion and assure participants that their responses will be used solely for research purposes and published anonymously. The “Personal Information” section assesses participants’ knowledge of software development and Continuous Integration, ensuring they have the necessary qualifications for the case study. The “Team Organization and Collaboration” section explores the organization’s CI environment, development processes, and tools used. Finally, the “Diagnosis of the Project’s CI Maturity” section investigates how the organization evaluates the CI maturity of its projects, the metrics employed, and overall satisfaction with the CI adoption.
3.2. Interviews
@@
We conduct semi-structured interviews (Dearnley, 2005) with the participants to obtain insights about monitoring CI practices of analyzed projects. We defined a set of initial questions, although they could vary depending on the participants’ responses during the interviews. The interview questions were used primarily to address the research questions RQ2 and RQ3 and are detailed in the case study protocol. We follow the guidance provided by (Hancock et al., 2021). We also follow the best practices before the Interview: 1) Identifying Key Participants, 2) Developing an Interview Guide/Protocol, 3) Ensuring a Private and Neutral Environment, 4) Recording the Interview Data, 5) Adhering to Legal and Ethical Requirements. Conducting the Interview we: 1) Obtaining Consent and Ensuring Anonymity, 2) Reviewing the Interview’s Purpose and Duration, 3) Listening Actively. We labeled each participant of case study to protect anonymity, Table 1 shows study participants.
Table 1. Case Study Participants
@@
3.3. Thematic Analysis of Interview Data
Following a similar approach to (Elazhary et al., 2021), we transcribed the interviews of all five participants. We assigned a label to participants’ statements in the interviews that sought to express, in a simple and summarized way, the participant’s feelings regarding monitoring CI practices. We grouped common themes and created a network graph to represent them. From there, we carried out an analysis to extract the main findings of these themes. We sent each participant’s final themes so they could review them and express whether they agreed with the extracted themes.
3.4. Study Replications
@@
https://zenodo.org/records/11106110. We provide the use case protocol (in English), the tool source code, as well as an example of the survey about the project’s CI maturity. Additionally, the values of CI practices throughout the case study and interview transcriptions are included anonymously (the latter three in Portuguese).
4. Results
4.1. Thematic Analysis
Figure 2 shows the main themes that emerge from the interviews about monitoring CI practices with the participants. Each theme summarizes an idea expressed by the participants during the interviews. Each theme has a number that represents the number of mentions expressed by participants. The green nodes are themes that have a positive aspect of monitoring CI practices, while the red nodes represent negative aspects pointed out by interview participants.
Figure 2. Interviews Themes
The primary positive theme that emerged when monitoring CI practices was the “Desire for Improvements” (28 mentions) which illustrates the participant’s desire to improve some of the CI practices within the project (i.e., the participant expresses the desire to improve certain practices in the future). For instance, participant P1 expressed this sentiment by stating: “We had already been doing this testing issue before, but it certainly reinforced the need to do it as quickly as possible.”. The second most cited theme was “Highlights the Problem” (23 mentions) indicating the discovery of previously unnoticed issues by the development team through the aid of the monitoring tool, participant P2 articulated this experience, saying: “In fact, I found it strange that the coverage dropped, I thought it had already been resolved, there is something there that is being executed and resetting the data.”. “Ease of Monitoring” was the third most cited theme (18 mentions), this theme reflects the ease that developers found during the case study to monitor CI practices using our tool, participant P4 commented: “No, no, it’s not expensive. It was very intuitive”. Monitoring CI practices “Reminders and Motivators” participant of case study (16 mentions) as the tool constantly highlighted issues to developers, motivating them to improve the CI process, participant P3 remarked: “… but I think that seeing a very prominent percentage gets attention and helps convince other developers to write tests”.
Monitoring CI practices also provides an “Overview of CI State” for the project (12 mentions) enabling developers to view all CI-related aspects in one place, as participant P2 described: “… Precisely the tool summarizes several things that perhaps we need to use several tools to get to this information.”. Additionally, monitoring fosters “Encouraging Communication” among team members (12 mentions), as mention by participant P5: “I think it’s interesting to know what the practices are like. Just the fact that we talked somewhat superficially about commits would already be a benefit.”. Moreover, it promotes “CI Improvements”(9 mentions) as participant P1 emphasized: “I am more focused on developing the features that are missing and working on test coverage, which has the most impact for us. I just uploaded some testing issues that will improve the coverage”. There were also other themes mentioned to a lesser extent, such as: “Impact on Project Quality” (4 mentions), “Encouraging Colaboration” (1 mentions),“Gave a Direction” (1 mentions), and “Encouraging Changes” (1 mentions).
The primary negative theme in monitoring CI practices is “Challenges in Monitoring”. It emerged from difficulties using the tool or suggestions for improvement indicated by participants during the case study. Most of these issues were related to integrating the tool with other CI tools already used by the team, aiming to facilitate the CI practices monitoring. For example, participant P3 said: “If there was something like a link in the Gitlab interface as Sonarqube has. I could click and be redirected to the tool … ”, participant P5 comment: “I don’t know if the tool already has some type of integration with communication tools, could be a integration with the internal chat that we use for communication … ”. Integration with other tools is beyond the initial scope of this study, but it expresses the need for CI tools like GitLab to integrate monitoring of CI practices into their CI pipelines.
The “Time Management Challenges” theme is mainly related to the difficulty of the team implementing improvements in CI practices due to a lack of time dedicated to this task. This is a common complaint, as developers usually have other tasks and cannot be dedicated exclusively to improving CI practices for their projects. Overall, the monitoring tool proved to be easy and intuitive. Monitoring CI practices has not proven costly for participants. Participant P4 comment: “I think the tool is very intuitive, very didactic …”. Participant P1 said: “It has no major impact on configuration, doesn’t generate extra work, and adds significant value to this aspect of CI practices.”. Participant P5 appreciated the alerts sent by the tool: “I noticed that you can receive it by email. I found the feedback from the emails interesting.”
4.2. RQ1: Were the developers satisfied with the project’s CI maturity levels?
@@
Organization A currently characterizes itself as operating at an INTERMEDIATE level of CI adoption. The classification at this level did not involve utilizing any specific CI practice. However, there is a recognized need for a strategic shift towards incorporating metrics into their CI process. The team believes that to start using metrics, it is necessary to carry out planning to define which CI practices would be appropriate for their context and how this would be implemented and used. When questioned about their satisfaction with the current CI maturity level, they said they were satisfied with the company’s CI level before using the monitoring suite tool. Furthermore, there is a perception that adopting automated database migration techniques, eliminating manual database upgrade routines, and incorporating fully automatic acceptance testing would significantly contribute to enhancing the project’s CI maturity. However, during the interviews, after starting to monitor CI practices, participant P1 expressed the need to focus more on automated testing: “In general, since the beginning of the project, I think so, but I am still not completely satisfied, mainly the coverage has not yet reached the reference value.”. Clearly, CI monitoring, with the constant sending of alerts, helped participants realize that there were additional areas for improvement beyond those mentioned in the initial survey. As participant P1 reported: “At the beginning, I didn’t even have a perspective on how it would be. The tool, like the study, also emphasizes the importance of providing an accurate idea of how the CI is performing, enabling individuals to understand the metrics and reference values”.
@@
Organization B identifies itself as operating at an INTERMEDIATE level of CI maturity. The classification at this level is not determined by specific CI practices but rather by the successful automation of critical processes: the build process, automated testing, automatic generation of Docker images, and the deployment of new versions in a dedicated test/approval environment. Despite the absence of metric-driven assessment, participants express contentment with the current state of CI within the company. Nevertheless, they envision further improvements to enhance CI effectiveness. Participants believe that greater code coverage pertaining to the performed tests, incorporating tools focused on code security, and utilizing Flyway to automate database version evolution would contribute significantly to elevating the overall maturity and efficacy of the project’s CI practices. However, CI monitoring helped them realize that there were still several aspects to be improved, as participant P2 described: “It’s something we’re still maturing, we started from scratch. I had little knowledge but never learned how to install, deploy, or build a pipeline. Then, one thing ended up helping the other, seeing these metrics and what needed improvement in our projects”.
@@
The Organization C characterizes its current level of CI maturity as a BASE level. This self-assessment lacks the incorporation of specific CI practices to precisely delineate the extent of adoption. Despite this, the case study participant expresses a notable dissatisfaction with the existing level of CI within the company, emphasizing that CI is still in its initial stages of implementation.
@@
Our results reveal that, excluding Organization C, which is still in the initial stages of CI adoption and conducts most processes manually, the other projects were satisfied with the CI maturity level before the case study. Although they did not utilize any metric-driven assessment to evaluate CI maturity, they considered automating builds and writing tests as the criteria for defining the CI maturity of projects. However, despite initial satisfaction, CI monitoring helped participants to notice several aspects of CI that needed to be improved in their projects.
4.3. RQ2: What benefits does CI monitoring bring?
@@
In Organization A, one of the most important benefits demonstrated by CI monitoring was its capacity to give an overview of the CI stage of a project. This feature is not usually directly implemented in CI services, as Participant P1 commented: “It brings precisely this visibility of those aspects that can be considered a problem. As I was talking about this coverage, we already have this information on the Sonarqube. But the other ones are new aspects that bring a more qualitative view of the CI practices we did not have in other tools … In GitLab, you have a view of the pipeline but do not have an aggregation like in this dashboard, in statistical terms.”. Monitoring CI practices also proved to be easy to use and without a high cost for Organization A. Participant P1 explained: “ … because it is easy to use, it brings value, even if the person does not use it very often. It has notifications so that the person is always there to know what the values are like. So, in short, it has no major impact on configuration, doesn’t generate extra work, and adds significant value to this aspect of CI practices”
@@
In Organization B, monitoring CI practices greatly sparked communication and the desire for improvements. Regarding communication, Participant P4 highlighted: “yes undoubtedly, we started to hold internal conversations about how to interact more efficiently, we also added revisions for each commit we applied, this was not a common practice…”. Participant P3 commented on the desire to improve: “The red images (in dashboard) of the build duration caught my attention. Maybe this sparked my interest in investigating and reducing the dependencies used” It also made developers notice overlooked problems. Participant P2 describes: “It helped to highlight… The coverage is evident; it was interesting because it showed that there was coverage data, and suddenly, it disappeared. I think it’s helping.”. Participant P3 also said: “I think it … pointing out that we are having problems with the code coverage metric. Regarding the build, we already felt that the build took a long time and that the VMs were slow, but not that much, more than 8 hours of build time. We have tests, tests that involve the instantiation of other modules in VMs, …., but almost 10h is a very high value.”
@@
In Organization C, Participant P5 stated that monitoring CI practices improve them by emphasizing: “Specifically, this concern of keeping builds always working, for example. The number of commits, worrying about always committing.”. It highlights the problems: “Having this awareness that the number of commits is important in the project. If there is no significant quantity, we may make commits with a lot of information at once. It helps to cover different aspects”. Monitoring CI practices also provides an overview of CI stage of a project, as commented by Participant P5: “It brings more confidence related to some CI needs, for sure. It provided a better view of this process.”. It also encourages communication: “We even had a conversation to discuss these metrics values.”.
@@
Some benefits were observed for the different projects, such as: 1) The reminders and motivation provided to developers were among the most beneficial points highlighted by participants. Participant P1 said: “… it is another thing to be constantly reminded that you have to do it here.”. Monitoring CI practices is not a priority feature in other tools as Participant P2 claims: “I found that the tool consolidates a lot of information in one place that sometimes we have to look for on several Sonarqube screens. It helped a lot. It synthesized it in one place, which I think is cool.”. Finally, the functionality of constantly sending alerts to participants was remembered as a very important aspect of monitoring. Participant P1 reported: “Now with the email alert that is sent weekly. It keeps this concern with metrics at the forefront of my mind. Otherwise, if you’re focused on other project problems, it’s easy to leave it aside”.
@@
Our findings show that the most important benefits demonstrated by CI monitoring were its capacity to give an overview of the CI stage, highlight CI-related problems, awaken a desire for improvements, improve the quality of CI adoption, and increase communication within the team. CI monitoring also proved to be not costly to apply.
4.4. RQ3: What are the challenges/disadvantages for using it on a daily basis?
@@
One of the biggest challenges reported in Organization A to use the monitoring of CI practices daily was not directly related to monitoring CI practices but to improving the CI maturity from the results displayed by the monitoring tool, mainly due to time issues. Monitoring CI practices helps highlight the project’s lack of testing, but there were obstacles to improving these practices, as revealed by Participant P1: “To generate a new project, we end up having to do a presentation during the sprint and sometimes I get a little worried about not having time to present what has been completed if I’m going to do all the tests, then we end up having to fit the tests into the free time we have in the sprint”
@@
When asked about the difficulties, Organization B also pointed out that the time dedicated to other development tasks was the main difficulty in monitoring and improving CI practices. Participant P2 said: “… We have very long builds with this build duration in another module, this one has a way to get around it, I just have not had the time yet.” and Participant P4: “… because the priority is the documentation that needs to be delivered by the end of the year”. The lack of integration with other CI tools was a problem also indicated as a difficulty in monitoring Organization B. Participant P3 said: “If there was something like a link in the Gitlab interface as Sonarqube has. I could click and be redirected to the tool … ” and Participant P4 mention: “maybe an idea, if you can find some way to add a plugin to Gitlab … ”
@@
In Organization C, relevant challenges in monitoring CI practices were also not reported. Legacy projects and the internal culture of the project were the main reasons that made improving CI practices, as shown by the tool, challenging. Participant P5 described the difficulty of adopting CI due to the characteristics of the projects, for example, in legacy systems: “It’s because it uses a legacy system, but it’s just a different way of working. The person who developed a large part of the project was a DBA. He worked directly with the tables, and in the end, it was complicated.” Participant P5 also suggested that better interaction with other tools could facilitate CI monitoring: “I don’t know if the tool already has some type of integration with communication tools, could be a integration with the internal chat that we use for communication … ”
@@
During the case study, we also observed additional issues that hinder the most appropriate CI adoption. In Organization A, the pressure for fast deliveries leads to underestimated deadlines, causing certain CI practices, such as writing automated tests, to be frequently postponed or ignored. Delayed testing may detect integration issues only when different parts of the system are combined, leading to significant delays and increased complexity in remediation. Organization B faces challenges due to a complex development environment, reflected in the instability of the CI environment, hindering any substantial improvement in the project’s CI maturity. Organization C needs to shift its project mindset to realize the full benefits of CI adoption. It was noted that other demands take priority, and developers are not actively encouraged to enhance CI maturity.
@@
Overall, there were no big challenges directly related to monitoring CI practices. The main challenges reported were related to improving CI practices due to a lack of time for developers who are required by other project tasks or due to the internal culture of the project. Some complaints were made about the tool’s usability and the lack of integration of the monitoring suite tool (graphical interface) with CI services. Pressure for fast deliveries, complexity and instability in the CI environment, and a lack of emphasis on the importance of CI were identified as problems during monitoring, hindering the improvement of CI maturity.
4.5. RQ4: Are developers interested in using the tool?

### v2

@@ III-A Survey
To gain an understanding of the project’s CI maturity before implementing the case study, we requested each developer to respond to a preliminary survey. This initial survey helped us understand the state of CI within each project prior to the case study, assess satisfaction with CI adoption, and identify any metrics used to measure CI. This preliminary insight facilitated a more comprehensive evaluation of the case study’s impact on the evolution of CI in each analyzed project. Furthermore, the survey responses allowed us to answer RQ1.
@@
The survey consisted of 15 questions, 9 of which were open-ended. To give respondents as much flexibility as possible, 9 questions were mandatory (6 of which were close-ended while 3 were open-ended), and 6 questions were optional (all open-ended). The survey was divided into four sections: 1) Introduction, 2) Personal Information, 3) Team Organization and Collaboration, and 4) Diagnosis of the project’s CI maturity.
@@
In the “Introduction” section of the survey, we provide context for the study, introduce the concept of CI, and outline the survey’s objectives. We also specify the estimated time required for completion and assure participants that their responses will be used solely for research purposes and published anonymously. The “Personal Information” section assesses participants’ knowledge of software development and Continuous Integration, ensuring they have the necessary qualifications for the case study. The “Team Organization and Collaboration” section explores the organization’s CI environment, development processes, and tools. Finally, the “Diagnosis of the Project’s CI Maturity” section investigates how the organization evaluates the CI maturity of its own projects, the metrics employed, and overall satisfaction with CI adoption.
III-B Interviews
@@
We conducted semi-structured interviews [24] to obtain insights into the monitoring practices regarding the CI of the analyzed projects. We defined a set of initial questions, although these could vary depending on the participants’ responses during the interviews. The interview questions were used primarily to address RQ2 and RQ3 and are detailed in the case study protocol. We followed the guidance provided by [25]. We also adopted best practices before the interviews, including: 1) identifying key participants, 2) developing an interview guide/protocol, 3) ensuring a private and neutral environment, 4) recording the interview data, and 5) adhering to legal and ethical requirements. When conducting the interviews, we: 1) obtained consent and ensured anonymity, 2) reviewed the interview’s purpose and duration, and 3) listened actively. We generated an alias for each participant to ensure anonymity. Table I shows the study participants.
TABLE I: Case Study Participants
@@
III-C Thematic Analysis of Interview Data
Following a similar approach to [9], we transcribed the interviews of all five participants. We assigned labels to participants’ statements that sought to express, in a simple and summarized manner, the participants’ perceptions regarding monitoring CI practices. We grouped common themes and created a network graph to represent them. From there, we conducted an analysis to extract the main findings from these themes. We shared each participant’s final themes with them so they could review and indicate whether they agreed with the extracted themes.
III-D Study Replications
@@
https://zenodo.org/records/15507762. We provide the use case protocol, the source code, as well as an example of the survey regarding the project’s CI maturity. The CI metrics’ values throughout the case study and interview transcriptions are included in an anonymous manner.
IV Results
IV-A RQ1: Are developers satisfied with the project’s CI maturity levels?
@@
Organization A currently characterizes itself as operating at an INTERMEDIATE level of CI adoption. This classification did not involve the use of any specific CI practice. However, there is a recognized need for a strategic shift toward incorporating such practices into their CI process. The team believes that to begin adopting practices, it is necessary to plan which CI practices are appropriate for their context and determine how they would be implemented and used. When asked about their satisfaction with the current CI maturity level, they stated they were satisfied with the company’s progress. Furthermore, there is a perception that adopting automated database migration techniques, eliminating manual database upgrade routines, and incorporating fully automated acceptance testing would significantly enhance the project’s CI maturity.
@@
Organization B identified itself as operating at an INTERMEDIATE level of CI maturity. The classification at this level is not determined by specific CI practices but rather by the successful automation of critical processes: the build process, automated testing, automatic generation of Docker images, and the deployment of new versions in a dedicated test/approval environment. Despite the absence of a metric-driven assessment, the participants expressed satisfaction with the current state of CI within the company. Nevertheless, they envisioned further improvements to enhance CI effectiveness. Participants believed that greater code coverage related to automated tests, incorporating tools focused on code security, and utilizing Flyway to automate database version evolution would contribute significantly to elevating the overall maturity and efficacy of the project’s CI practices.
@@
Organization C considered its current level of CI maturity as BASE level. This level lacks the incorporation of specific CI practices to precisely delineate the extent of adoption. Despite this, the case study participant expresses a notable dissatisfaction with the existing level of CI within the company, emphasizing that CI is still in its initial stages of implementation.
@@
Our results reveal that, except for Organization C — which is still in the early stages of CI adoption and performs most processes manually — the other projects were satisfied with their CI maturity level prior to the case study. Although they did not use any metric-driven assessment to evaluate CI maturity, they considered build automation and test writing as key criteria for defining a project’s CI maturity.
IV-B RQ2: What benefits does CI monitoring bring?
@@
In Organization A, one of the most important benefits demonstrated by CI monitoring was its capacity to give an overview of a project’s CI stage. This feature is not usually directly implemented in CI services, as Participant P1 commented: “It brings precisely the visibility of those aspects that can be considered a problem. As I was talking about this coverage, we already have this information on the Sonarqube. But the other ones are new aspects that bring a more qualitative view of the CI practices we did not have in other tools … In GitLab, you have a view of the pipeline but do not have an aggregation like in this dashboard, in statistical terms.”. Monitoring CI practices also proved to be easy to use and low-cost for Organization A. Participant P1 explained: “ … because it is easy to use, it brings value, even if the developer does not use it very often. It has notifications so that developers are always there to know what the values are like. So, in short, it has no major impact on configuration, does not generate extra work, and adds significant value to this aspect of CI practices”
@@
In Organization B, monitoring CI practices significantly stimulated communication and a desire for improvements. Regarding communication, Participant P4 highlighted: “yes undoubtedly, we started to hold internal conversations about how to interact more efficiently, we also added revisions for each commit we applied, this was not a common practice…”. Participant P3 commented on the desire to improve: “The red images (in dashboard) of the build duration caught my attention. Maybe this sparked my interest in investigating and reducing the dependencies used”. Monitoring also helped developers notice previously overlooked problems. Participant P2 described: “It helped to highlight… The coverage is evident; it was interesting because it showed that there was coverage data, and then suddenly it disappeared. I think it’s helping.”. Participant P3 also said: “I think it … pointing out that we are having problems with the code coverage metric. Regarding the build, we already felt that the build took a long time and that the VMs were slow, but not that much, more than 8 hours of build time. We have tests that involve the instantiation of other modules in VMs, …., but almost 10 hours is a very high value”.
@@
In Organization C, Participant P5 stated that monitoring CI practices help improve them by emphasizing: “Specifically, this concern of keeping builds always working, for example. The number of commits, worrying about always committing.”. It also highlights problems, as noted by P5: “Having this awareness that the number of commits is important in the project. If there is no significant quantity, we may make commits with a lot of information at once. It helps to cover different aspects”. Monitoring CI practices also provides an overview of CI stage of a project, as commented by Participant P5: “It brings more confidence related to some CI needs, for sure. It provided a better view of this process.”. Finally, it encourages communication: “We even had a conversation to discuss these metrics values.”.
@@
Some benefits were observed across the different projects. One of the most beneficial points highlighted by participants was the reminders and motivation provided to developers. Participant P1 said: “… it is another thing to be constantly reminded that you have to do it here.”. Monitoring CI practices is not a priority feature in other tools, as Participant P2 stated: “I found that the tool consolidates a lot of information in one place that sometimes we have to look for on several Sonarqube screens. It helped a lot. It synthesized it in one place, which I think is cool.”. The functionality of sending regular alerts to participants was also mentioned as a very important aspect of monitoring. Participant P1 reported: “Now with the email alert that is sent weekly. It keeps this concern with metrics at the forefront of my mind. Otherwise, if you’re focused on other project problems, it’s easy to leave it aside”.
@@
Our findings show that the most important benefits demonstrated by CI monitoring were its ability to provide an overview of the CI stage, highlight CI-related problems, awaken a desire for improvements, improve the quality of CI adoption, and increase communication within the team. CI monitoring also proved to be cost-effective.
IV-C RQ3: What are the challenges/disadvantages of monitoring CI practices on a daily basis?
@@
One of the biggest challenges reported in Organization A in using CI practice monitoring daily was not directly related to the monitoring itself, but to acting on the results it provided. Specifically, improving CI maturity, mainly due to time constraints. Monitoring helped highlight the project’s lack of testing, but there were obstacles to addressing these issues, as revealed by Participant P1: “… we end up having to prepare a presentation during the sprint and sometimes I get a little worried about not having time to present what has been completed. If I am going to implement several automated tests, then we end up having to fit the tests into the free time we have in the sprint”.
@@
When asked about difficulties, Organization B also pointed out that the time dedicated to other development tasks was the main challenge in monitoring and improving CI practices. Participant P2 said: “… We have very long builds with this build duration in another module, this one has a way to get around it, I just have not had the time yet.” and Participant P4: “… because the priority is the documentation that needs to be delivered by the end of the year”. The lack of integration with other CI tools was another problem identified as a difficulty in monitoring Organization B. Participant P3 said: “If there was something like a link in the Gitlab interface as Sonarqube has. I could click and be redirected to the tool … ” and Participant P4 mentioned: “maybe an idea, if you can find some way to add a plugin to Gitlab … ”
@@
In Organization C, no relevant challenges in monitoring CI practices were reported. Legacy projects and the internal culture of the project were the main reasons that made improving CI practices, as shown by the tool, challenging. Participant P5 described the difficulty of adopting CI due to the characteristics of the projects, especially legacy systems: “It’s because it uses a legacy system, but it’s just a different way of working. The person who developed a large part of the project was a DBA. He worked directly with the tables, and in the end, it was complicated.” Participant P5 also suggested that better interaction with other tools could facilitate CI monitoring: “I don’t know if the tool already has some type of integration with communication tools, could be a integration with the internal chat that we use for communication … ”
@@
During the case study, we also observed additional issues that hinder the appropriate adoption of CI. In Organization A, the pressure for fast deliveries leads to underestimated deadlines, causing certain CI practices, such as writing automated tests, to be frequently postponed or ignored. Delayed testing may detect integration issues only when different parts of the system are combined, resulting to significant delays and increased complexity in remediation. Organization B faces challenges due to a complex development environment, reflected in the instability of the CI environment, which hinders any substantial improvement in the project’s CI maturity. Organization C needs to shift its project mindset to realize the full benefits of CI adoption. It was noted that other demands take priority, and developers are not actively encouraged to enhance CI maturity.
@@
Overall, there were no major challenges directly related to monitoring CI practices. The main challenges reported concerned improving CI practices due to developers’ lack of time caused by other project tasks or the internal culture of the project. Some complaints were made about the tool’s usability and the lack of integration between the monitoring suite (graphical interface) and CI services. Pressure for fast deliveries, complexity and instability in the CI environment, and a lack of emphasis on the importance of CI were identified as problems during monitoring, hindering the improvement of CI maturity.
IV-D RQ4: Are developers interested in using our metrics suite tool?